OpenStudy (anonymous):

2x-2 y+24 5y-8 x+2 Find values of x and y for the paralellogram

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks!... but what exactly are the x and y values?

OpenStudy (anonymous):

do u know that opposite sides of parallelogram are equal use this property to first from equations in x and y and then furthe use it to solve and get the x and y

OpenStudy (anonymous):

if what ayushprasad has drawn for u is correct then ur equation stand as 2x-2=x+2 and 5y-8 =y+24 solve the above equations and check urself that u get x=4 and y=8

OpenStudy (anonymous):

2(4)-2=4+2 5(8)-8=8+24 It's correct!!! thanks @matricked !!

OpenStudy (anonymous):

welcome dear